OpenStudy (anonymous):

Given sin theta=-1. theta= ? a. 180 degrees b. 0 degrees c. 270 degrees d. 90 degrees

OpenStudy (anonymous):

When is sine -1 on the unit circle? Since since is the y-coordinate, a -1 would be between quadrant 3 and 4 on the y-axis, so at 270 degrees

OpenStudy (anonymous):

\[\sin \theta = -1 \implies \sin \theta = -\sin(90) \implies \sin(-90) \implies \sin(2 \pi - 90) = \sin(270)\] \[\sin \theta = \sin(270)\] \[\theta = 270\]

OpenStudy (anonymous):

Thanks guys